OpenStudy (anonymous):

What are the steps to solve this problem? 0.000027 divided by 0.000009 =

OpenStudy (azureilai):

you can plug the numbers into a calculator, or use scientific notation

OpenStudy (anonymous):

multiply both sides by 1,000,000 (move the decimal over to the right 6 places)

OpenStudy (anonymous):

\[0.000027=\frac{ 27 }{1000000 },.000009=\frac{ 9 }{1000000 }\] 0.000027/0.000009 \[=\frac{ 27 }{ 1000000 }*\frac{ 1000000 }{9 }=3\] or \[\frac{ 27*10^{-6} }{9*10^{-6} }=3\]